# INSIGHT: End-to-End Neuro-Symbolic Visual Reinforcement Learning with Language Explanations

### v2

@@ Abstract
Neuro-symbolic reinforcement learning (NS-RL) has emerged as a promising paradigm for explainable decision-making, characterized by the interpretability of symbolic policies. NS-RL entails structured state representations for tasks with visual observations, but previous methods are unable to refine the structured states with rewards due to a lack of efficiency. Accessibility also remains to be an issue, as extensive domain knowledge is required to interpret symbolic policies. In this paper, we present a framework for learning structured states and symbolic policies jointly, whose key idea is to distill vision foundation models into a scalable perception module and refine it during policy learning. Moreover, we design a pipeline to generate language explanations for policies and decisions using large language models. In experiments on nine Atari tasks, we verify the efficacy of our approach, and we also present explanations for policies and decisions.
@@ 1 Introduction
Recent years have witnessed remarkable progress of deep reinforcement learning (RL) (Agarwal et al., 2021; Wurman et al., 2022; Degrave et al., 2022). However, deep RL still faces limitations in sensitive domains due to its opaque nature (Milani et al., 2022). For example, the opacity hinders diagnosing or rectifying policies that fails to generalize, which may be caused by exploiting peripheral information (Delfosse et al., 2024). Neuro-symbolic reinforcement learning (NS-RL) is promising for overcoming this limitation (Verma et al., 2018; Coppens et al., 2019). It uses structured states and parameterizes policies with concise expressions, thereby guaranteeing interpretability via clear semantics of states and policies.
@@
Meanwhile, symbolic policies can be intricate for general users, albeit being intrinsically interpretable. For instance, to interpret logical policies (Delfosse et al., 2023b) one has to be familiar with first-order logic, and in the case of programmatic policies (Qiu & Zhu, 2022) one has to learn the corresponding grammars. Such a lack of accessibility can be an obstacle for agents to gain the trust of the general public. Nevertheless, there is a lack effort in the literature of NS-RL to explain policies for non-expert users.
In this paper, we present a framework called INSIGHT (INSIGHT) to address these issues. As illustrated in Fig. 1, INSIGHT can learn the object coordinates and symbolic policies simultaneously and explain policies and specific decisions in natural language. The key idea of INSIGHT is to overcomes the efficiency drawback of previous methods by distilling vision foundation models into a scalable perception module. Accompanied by the equation learner (EQL) (Sahoo et al., 2018) for representing policies, both the perception module and symbolic policies can be learned from rewards in an end-to-end fashion. Moreover, we develop a pipeline for explaining policies and decisions in natural language using large language models (LLMs), which reduces the cognitive load on users to understand symbolic policies learned by INSIGHT. Through a step called concept grounding, LLMs are instructed to associate quantities in symbolic policies with their semantics for the task of interest. Then, as illustrated in Fig. 1, LLMs are prompted for explaining the decision-making patterns of a symbolic policy and specific decisions made by the policy.
We verify the efficacy of INSIGHT and its model design with extensive experiments on nine Atari games. INSIGHT outperforms all existing approaches for NS-RL (CGP (Wilson et al., 2018), Diffses (Yuan et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023b)). We show that its empirical performance can be attributed to the refinement of structured states with reward information and our approach for learning symbolic policies. We also present examples of policy interpretations and decision explanations. In summary, our contributions are three-fold.
@@
    We develop a pipeline for explaining symbolic policies in natural languages.
@@
    We demonstrate the efficacy of our framework on nine Atari games and showcase language explanations.
@@ 2 Related work
Approaches for explainable RL can be categorized into post hoc and intrinsic methods. The former generates explanations using predefined templates (Ehsan et al., 2018; Wang et al., 2019; Hayes & Shah, 2017) or saliency maps (Greydanus et al., 2018), which can be subjective and unreliable (Dazeley et al., 2023). The latter relies on interpretable policy classes, such decision trees (Topin et al., 2021; Zhang et al., 2021), logical expressions (Dazeley et al., 2023; Delfosse et al., 2023b, 2024), and mathematical expressions (Zheng et al., 2022; Landajuela et al., 2021), for better transparency. Meanwhile, we argue that intrinsic transparency does not guarantee accessibility. While logic-based approaches are better for characterizing relations, they can be unfriendly to users due to their prerequisite knowledge. Recently, there is a growing interest in using LLMs to explain machine learning models (Kroeger et al., 2023; Tennenholtz et al., 2023; Singh et al., 2023; Bills et al., 2023; Zhang et al., 2023). INSIGHT is the first NS-RL method that improves accessibility by using LLMs to generate explanations automatically.
Meanwhile, prior NS-RL approaches (Verma et al., 2018; Coppens et al., 2019; Verma et al., 2019; Landajuela et al., 2021) focus on tasks with low-dimensional observations that have clear semantics. For tasks with visual input, they either rely on ground truth (Delfosse et al., 2023b) or human-defined primitives (Wilson et al., 2018; Lyu et al., 2019) for state representations. As an exception, Zheng et al. (2022) proposed to extract structured states from pre-trained SPACE models, yet the representations are not refined with reward signals and lead to significant performance degradation. INSIGHT is the first NS-RL approach that learns the structured states from both visual and reward signals.
@@ 3.2 Learning Symbolic Policies
EQL  We represent symbolic policies with the EQL network (Martius & Lampert, 2016; Sahoo et al., 2018). Given an input vector $\mathbf{h}^{(k-1)}\in\mathbb{R}^{d_{k-1}}$, its $k$th layer first applies a transformation $\mathbf{g}^{(k)}=\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}+\mathbf{b}^{(k)}$, where $\mathbf{W}^{(k)}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ and $\mathbf{b}^{(k)}\in\mathbb{R}^{d_{k}}$ are learnable parameters. $d_{k-1}$ and $d_{k}$ are the dimension of $\mathbf{h}^{(k-1)}$ and $\mathbf{g}^{(k)}$.
@@
The final ingredient is sparsity regularization (Martius & Lampert, 2016) required for deriving concise expressions, which imposes the following regularization to the parameters of a EQL network to avoid the singularity in the gradient as the weights go to 0.
@@
Table 1: INSIGHT matches the performance of neural agents on all Atari tasks. Performance of the proposed INSIGHT, Neural (a neural agent that has the same network architecture as INSIGHT), and existing NS-RL approaches (CGP (Wilson et al., 2018), Diffses (Zheng et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2023b)) on nine Atari tasks. INSIGHT outperforms all NS-RL baselines and matches the performance of Neural for all nine tasks.
@@
    We used the results reported by Delfosse et al. (2023b) for Freeway and obtained results for Pong, Enduro, SpaceInvaders, Seaquest, and Breakout using its released code.
@@ 4 Experiments
This section evaluates the efficacy of INSIGHT for learning symbolic policy and structured state representations. Sec. 4.2 reports the task performance of INSIGHT and validates our design choices, and Sec. 4.3 presents language explanations for learned policies and specific decisions.
@@ 4.1 Experimental Setup
Tasks  We consider the online RL setting and nine Atari tasks: Pong, BeamRider, Enduro, Qbert, SpaceInvaders, Seaquest, Breakout, Freeway, and MsPacman, which widely used in prior works (Delfosse et al., 2023b; Landajuela et al., 2021; Zheng et al., 2022). They range from simple motion control (e.g. Pong) to complex decision-making (e.g. Seaquest), and cover issues such as clean background (e.g. SpaceInvaders) vs noisy background (e.g. BeamRider) and constant object (e.g. Pong) vs varying objects (e.g. Enduro). So with them we can examine INSIGHT comprehensively. In the meantime, results on more realistic tasks are also important for validating NS-RL algorithms, so we also include results on MetaDrive (Li et al., 2022), a challenging environment for autonomous driving.
@@
Baselines  For task performance, INSIGHT is compared with Neural, DSP, Diffses, NUDGE, and CGP. Neural is a deep RL alternative for INSIGHT that uses the same network architecture but does not learn from the frame-symbol datasets. The remaining four are approaches for NS-RL. DSP (Landajuela et al., 2021) searches for symbolic policies from low-dimensional input using a recurrent neural network, while Diffses (Zheng et al., 2022) uses frozen SPACE models to extract states and search for symbolic policies using genetic programming. NUDGE (Delfosse et al., 2023b) relies on the ground-truth structured states and represent policies with first-order differentiable logic. Lastly, CGP (Wilson et al., 2018) leverages cartesian genetic programming for learning policies from images.
@@
Implementation Details  We use the standard pre-processing protocol for Atari tasks, which includes resizing, gray scaling, and frame stacking. As for hyper-parameters, we use 1e-3 for $\lambda_{\text{reg}}$, 2 for $\lambda_{\text{cnn}}$. We use the OpenAI’s GPT-4 model (Bubeck et al., 2023) when generating explanations. All the implementation details are provided in Sec. B.1.
@@ 4.2 Results
Figure 5: Examples for language explanations for Pong. Readers may refer to Appx. D for full prompts. Left: interpretations for a learned policy. The interpretations identify influential input variables and summarize triggering patterns of actions. Right: explanations for an action taken at a state. The four images located at the bottom illustrate the state. The motion of the ball and the opponent’s paddle are deduced from input variables, which are used for supporting explanations of actions.
@@ 4.3 Language Explanations
The right part of Fig. 5 showcases decision explanations. While the movements of the ball is correctly deduced, there is a mistake for the agent’s paddle, possibly due to fluctuations in the coordinate predictions. The generated explanations reveals that the decision is more sensitive to the position of the opponent’s paddle than the ball, suggesting that the agent is exploiting the opponent’s fixed policy. Similar phenomena are also observed in recent NS-RL works (Delfosse et al., 2024), which highlights the importance of inspecting learned policies manually. Nevertheless, INSIGHT can generate credible explanations by associating interpretations with facts.
@@ 5 Conclusions
A limitation of INSIGHT is that EQL networks cannot express logical operations required by some reasoning tasks. Furthermore, quantitative evaluation of the explanations is also an interesting future topic.
@@ Acknowledgments
C. Fang was supported by National Science and Technology Major Project (2022ZD0114902), the NSF China (No. 62376008). The authors would like to thank four anonymous reviews for constructive feedback.
@@
The goal of this paper is to improve the transparency of RL without sacrificing task performance, which paves the road towards trustworthy RL applications. From a societal point of view, we aim to reduce the barrier between general public and RL models, thus advancing the broad application of RL agents.
@@ C.1 Comprehensive Evaluation of Representations
Distilled Structured Representations Significantly Enhance Performance These methods were trained for only one million steps due to the poor efficiency of SPACE-Neural. Similar experiment setup has been utilized by Delfosse et al. (2023c). We also report results for three million steps in Tab. A6, by which point the performances on Pong and Freeway had stabilized. Note that compared to slot attention or the SPACE model, our perception module leads to better performance. This is because our image-symbol dataset provides direct supervision for the location of objects. On the contrary, unsupervised methods are trained to reconstruct whole images and thus less focused on individual objects.
@@ C.6 Accuracy Evaluation of the FastSAM and DeAot Method
Our Approach Competes with MOC Solely Through Zero-Shot Generalization To evaluate the reliability of our dataset generation method, we further evaluate the accuracy of bounding boxes in the OCAtari dataset  (Delfosse et al., 2023a). The table below shows the F-score from the Pong, SpaceInvaders and MsPacman environments within the OCAtari dataset, compared with the results of SPACE+MOC, as reported in  (Delfosse et al., 2023d):

### v3

# End-to-End Neuro-Symbolic Reinforcement Learning with Textual Explanations

@@ Abstract
Neuro-symbolic reinforcement learning (NS-RL) has emerged as a promising paradigm for explainable decision-making, characterized by the interpretability of symbolic policies. NS-RL entails structured state representations for tasks with visual observations, but previous methods cannot refine the structured states with rewards due to a lack of efficiency. Accessibility also remains an issue, as extensive domain knowledge is required to interpret symbolic policies. In this paper, we present a neuro-symbolic framework for jointly learning structured states and symbolic policies, whose key idea is to distill the vision foundation model into an efficient perception module and refine it during policy learning. Moreover, we design a pipeline to prompt GPT-4 to generate textual explanations for the learned policies and decisions, significantly reducing users’ cognitive load to understand the symbolic policies. We verify the efficacy of our approach on nine Atari tasks and present GPT-generated explanations for policies and decisions.
@@ 1 Introduction
Recent years have witnessed remarkable progress of deep reinforcement learning (RL) (Agarwal et al., 2021; Wurman et al., 2022; Degrave et al., 2022). However, deep RL still faces limitations in sensitive domains due to its opaque nature (Milani et al., 2022). For example, the opacity hinders diagnosing or rectifying policies that fails to generalize, which may be caused by exploiting peripheral information (Delfosse et al., 2024b). Neuro-symbolic reinforcement learning (NS-RL) is promising for overcoming this limitation (Verma et al., 2018; Coppens et al., 2019). It uses structured states and parameterizes policies with concise expressions, thereby guaranteeing interpretability via clear semantics of states and policies.
@@
Meanwhile, symbolic policies can be intricate for general users, albeit being intrinsically interpretable. For instance, to interpret logical policies (Delfosse et al., 2024a) one has to be familiar with first-order logic, and in the case of programmatic policies (Qiu & Zhu, 2022) one has to learn the corresponding grammars. Such a lack of accessibility can be an obstacle for agents to gain the trust of the general public. Nevertheless, there is a lack effort in the literature of NS-RL to explain policies for non-expert users.
To address these issues, we present INSIGHT, an interpretable neuro-symbolic framework for visual reinforcement learning. As illustrated in Fig. 1, INSIGHT can learn the object coordinates and symbolic policies simultaneously and explain policies and specific decisions in natural language. The key idea of INSIGHT is to overcomes the efficiency drawback of previous methods by distilling vision foundation models into a scalable perception module. Accompanied by the equation learner (EQL) (Sahoo et al., 2018) for representing policies, both the perception module and symbolic policies can be learned from rewards in an end-to-end fashion. Moreover, we develop a pipeline for explaining policies and decisions in natural language via GPT-4, which reduces the cognitive load on users to understand symbolic policies learned by INSIGHT. Through a step called concept grounding, LLMs are instructed to associate quantities in symbolic policies with their semantics for the task of interest. Then, as illustrated in Fig. 1, LLMs are prompted for explaining the decision-making patterns of a symbolic policy and specific decisions made by the policy.
We verify the efficacy of INSIGHT and its model design with extensive experiments on nine Atari games. INSIGHT outperforms all existing approaches for NS-RL (CGP (Wilson et al., 2018), Diffses (Yuan et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2024a)). We show that its empirical performance can be attributed to the refinement of structured states with reward information and our approach for learning symbolic policies. We also present examples of policy interpretations and decision explanations.
In summary, our contributions are three-fold.
@@
    We develop a pipeline to prompt GPT-4 to explain the learned policies and decisions in natural language.
@@
    We demonstrate the efficacy of our framework on nine Atari games and showcase textual explanations.
@@ 2 Related work
Approaches for explainable RL can be categorized into post hoc and intrinsic methods. The former generates explanations using predefined templates (Ehsan et al., 2018; Wang et al., 2019; Hayes & Shah, 2017) or saliency maps (Greydanus et al., 2018), which can be subjective and unreliable (Dazeley et al., 2023). The latter relies on interpretable policy classes, such decision trees (Topin et al., 2021; Zhang et al., 2021), logical expressions (Dazeley et al., 2023; Delfosse et al., 2024a, b), and mathematical expressions (Zheng et al., 2022; Landajuela et al., 2021), for better transparency. Meanwhile, we argue that intrinsic transparency does not guarantee accessibility. While logic-based approaches are better for characterizing relations, they can be unfriendly to users due to their prerequisite knowledge. Recently, there is a growing interest in using LLMs to explain machine learning models (Kroeger et al., 2023; Tennenholtz et al., 2023; Singh et al., 2023; Bills et al., 2023; Zhang et al., 2023). INSIGHT is the first NS-RL method that improves accessibility by using LLMs to generate explanations automatically.
While neuro-symbolic approaches (Manhaeve et al., 2018; Li et al., 2020, 2024) have better interpretability, prior NS-RL approaches (Verma et al., 2018; Coppens et al., 2019; Verma et al., 2019; Landajuela et al., 2021) focus on tasks with low-dimensional observations that have clear semantics. For tasks with visual input, they either rely on ground truth (Delfosse et al., 2024a) or human-defined primitives (Wilson et al., 2018; Lyu et al., 2019) for state representations. As an exception, Zheng et al. (2022) proposed to extract structured states from pre-trained SPACE models, yet the representations are not refined with reward signals and lead to significant performance degradation. INSIGHT is the first NS-RL approach that learns the structured states from both visual and reward signals.
@@ 3.2 Learning Symbolic Policies
EQL  We represent symbolic policies with the EQL network (Martius & Lampert, 2017; Sahoo et al., 2018). Given an input vector $\mathbf{h}^{(k-1)}\in\mathbb{R}^{d_{k-1}}$, its $k$th layer first applies a transformation $\mathbf{g}^{(k)}=\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}+\mathbf{b}^{(k)}$, where $\mathbf{W}^{(k)}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ and $\mathbf{b}^{(k)}\in\mathbb{R}^{d_{k}}$ are learnable parameters. $d_{k-1}$ and $d_{k}$ are the dimension of $\mathbf{h}^{(k-1)}$ and $\mathbf{g}^{(k)}$.
@@
The final ingredient is sparsity regularization (Martius & Lampert, 2017) required for deriving concise expressions, which imposes the following regularization to the parameters of a EQL network to avoid the singularity in the gradient as the weights go to 0.
@@
Table 1: INSIGHT matches the performance of neural agents on all Atari tasks. Performance of the proposed INSIGHT, Neural (a neural agent that has the same network architecture as INSIGHT), and existing NS-RL approaches (CGP (Wilson et al., 2018), Diffses (Zheng et al., 2022), DSP (Landajuela et al., 2021), and NUDGE (Delfosse et al., 2024a)) on nine Atari tasks. INSIGHT outperforms all NS-RL baselines and matches the performance of Neural for all nine tasks.
@@
    We used the results reported by Delfosse et al. (2024a) for Freeway and obtained results for Pong, Enduro, SpaceInvaders, Seaquest, and Breakout using its released code.
@@ 4 Experiments
This section evaluates the efficacy of INSIGHT for learning symbolic policy and structured state representations. Sec. 4.2 reports the task performance of INSIGHT and validates our design choices, and Sec. 4.3 presents textual explanations for learned policies and specific decisions.
@@ 4.1 Experimental Setup
Tasks  We consider the online RL setting and nine Atari tasks: Pong, BeamRider, Enduro, Qbert, SpaceInvaders, Seaquest, Breakout, Freeway, and MsPacman, which widely used in prior works (Delfosse et al., 2024a; Landajuela et al., 2021; Zheng et al., 2022). They range from simple motion control (e.g. Pong) to complex decision-making (e.g. Seaquest), and cover issues such as clean background (e.g. SpaceInvaders) vs noisy background (e.g. BeamRider) and constant object (e.g. Pong) vs varying objects (e.g. Enduro). So with them we can examine INSIGHT comprehensively. In the meantime, results on more realistic tasks are also important for validating NS-RL algorithms, so we also include results on MetaDrive (Li et al., 2022), a challenging environment for autonomous driving.
@@
Baselines  For task performance, INSIGHT is compared with Neural, DSP, Diffses, NUDGE, and CGP. Neural is a deep RL alternative for INSIGHT that uses the same network architecture but does not learn from the frame-symbol datasets. The remaining four are approaches for NS-RL. DSP (Landajuela et al., 2021) searches for symbolic policies from low-dimensional input using a recurrent neural network, while Diffses (Zheng et al., 2022) uses frozen SPACE models to extract states and search for symbolic policies using genetic programming. NUDGE (Delfosse et al., 2024a) relies on the ground-truth structured states and represent policies with first-order differentiable logic. Lastly, CGP (Wilson et al., 2018) leverages cartesian genetic programming for learning policies from images.
@@
Implementation Details  We use the standard pre-processing protocol for Atari tasks, which includes resizing, gray scaling, and frame stacking. As for hyper-parameters, we use $0.001$ for $\lambda_{\text{reg}}$, 2 for $\lambda_{\text{cnn}}$. We use the OpenAI’s GPT-4 model (Bubeck et al., 2023) when generating explanations. All the implementation details are provided in Sec. B.1.
@@ 4.2 Quantitative Results
Figure 5: Examples for textual explanations for Pong. Readers may refer to Appx. D for full prompts. Left: interpretations for a learned policy. The interpretations identify influential input variables and summarize triggering patterns of actions. Right: explanations for an action taken at a state. The four images located at the bottom illustrate the state. The motion of the ball and the opponent’s paddle are deduced from input variables, which are used for supporting explanations of actions.
@@ 4.3 Textual Explanations
The right part of Fig. 5 showcases decision explanations. While the movements of the ball is correctly deduced, there is a mistake for the agent’s paddle, possibly due to fluctuations in the coordinate predictions. The generated explanations reveals that the decision is more sensitive to the position of the opponent’s paddle than the ball, suggesting that the agent is exploiting the opponent’s fixed policy. Similar phenomena are also observed in recent NS-RL works (Delfosse et al., 2024b), which highlights the importance of inspecting learned policies manually. Nevertheless, INSIGHT can generate credible explanations by associating interpretations with facts.
@@ 5 Conclusions
Currently, the EQL network in INSIGHT cannot express logical operations required by some reasoning tasks. Furthermore, quantitative evaluation of the explanations is also an interesting future topic.
@@ Acknowledgments
C. Fang was supported by the National Science and Technology Major Project (2022ZD0114902), the NSF China (No. 62376008). The authors would like to thank four anonymous reviewers for constructive feedback.
@@
This paper aims to improve RL’s transparency without sacrificing task performance, which paves the road towards trustworthy RL applications. From a societal point of view, we aim to reduce the barrier between the general public and RL models, thus advancing the broad application of RL agents.
@@ C.1 Comprehensive Evaluation of Representations
Distilled Structured Representations Significantly Enhance Performance These methods were trained for only one million steps due to the poor efficiency of SPACE-Neural. Similar experiment setup has been utilized by Delfosse et al. (2023b). We also report results for three million steps in Tab. A6, by which point the performances on Pong and Freeway had stabilized. Note that compared to slot attention or the SPACE model, our perception module leads to better performance. This is because our image-symbol dataset provides direct supervision for the location of objects. On the contrary, unsupervised methods are trained to reconstruct whole images and thus less focused on individual objects.
@@ C.6 Accuracy Evaluation of the FastSAM and DeAot Method
Our Approach Competes with MOC Solely Through Zero-Shot Generalization To evaluate the reliability of our dataset generation method, we further evaluate the accuracy of bounding boxes in the OCAtari dataset  (Delfosse et al., 2023a). The table below shows the F-score from the Pong, SpaceInvaders and MsPacman environments within the OCAtari dataset, compared with the results of SPACE+MOC, as reported in  (Delfosse et al., 2023c):